# Radiation forces and torques in optics and acoustics

### v1

@@ IV.2.1 Chiral particles
Another important case, which appears only in electromagnetism, is a cross-coupling between the electric and magnetic induced dipoles and the electric and magnetic incident fields. This coupling occurs for isotropic chiral particles which lack the mirror and space-inversion (P) symmetries. Such particles typically consist of chiral micro-elements (e.g., molecules) with one prevailing handedness and random orientations (to ensure macroscopic isotropy). In this case, the relations for the induced electric and magnetic dipoles become [57]
where the parameter quantifies the chirality of the particle [344, 565].
@@
Substituting Eq. (31) into general equations of Table 2, and using definitions (3)–(6) we derive the additional absorption rates, forces, and torques induced by the chirality, and linear in [141, 227, 239]:
@@
These equations include both the main (‘mixed’) and recoil terms, and require some explanations.
@@
First, the chirality-dependent absorption rate (IV.2.1) describes the chiral dichroism, which plays a crucial role in optical sensing of chiral molecules [57, 530, 531, 260, 148, 97]. The main term there is proportional to the electromagnetic helicity density (6), which can also be associated with the “electromagnetic chirality” [530, 98, 123]. The additional recoil term involves a natural complexification of the helicity (6): , which was discussed in [97, 294, 403].
@@
Second, the chiral force (IV.2.1) contains three terms strongly resembling the gradient, radiation-pressure, and recoil forces in Eqs. (22) and (24). In this manner, the chiral gradient force is determined by the gradient of the helicity density in the incident field, whereas the chiral radiation-pressure force is proportional to a quantity which can be called “chiral momentum” density [128, 97, 556]:
@@
where we used Maxwell’s equations. The chiral momentum density characterizes the difference between the momenta of the right-hand and left-hand circularly-polarized photons; for the helicity eigenstates , , it equals . Importantly, the helicity-gradient and chiral radiation-pressure forces in Eq. (IV.2.1) have been successfully employed for the optical sorting (separation) of chiral particles [128, 97, 124, 539, 541, 565, 307] (see Section V.5).
@@
Third, we wrote the chiral optical torque (34) using the chiral absorption cross-sections [239]
@@
Akin to the isotropic-particle case, Eqs. (26), the chiral torque is determined by absorption and vanishes for lossless particles. Notably, this conclusion is valid for isotropic chiral spheres of any size [141], which is related to the optical theorem for chiral particles [72]. The chiral torque (34) provides another mechanism of optical torque, Fig. 9(d). Even a linearly-polarized incident plane wave produces such torque, because an absorptive chiral particle scatters this plane wave into a field with non-zero angular momentum. To the best of our knowledge, this kind of torque has not been measured experimentally yet, but it should be feasible for experiments with chiral metallic nano-structures [346, 225, 596].
@@
In addition to the chiral effects linear in , there are small corrections to the isotropic-particle force and torque, Eqs. (24) and (26): and . However, these corrections cannot discriminate between opposite chiralities of the particle (because of the invariance), and only provide small, mostly-negligible corrections to the magnitude of the isotropic recoil force and torque.
@@
Note that some works group the chiral and non-chiral force terms differently, so that additional ‘vortex’ and ‘spin-curl’ force terms appear [391, 565, 333, 139]. This occurs because of the absence of the canonical momentum and chiral momentum (35) in such approaches. These quantities provide a clear intuitive framework described here. The equivalence of different approaches is examined in detail in [239].
@@ IV.2.2 Magnetoelectric coupling
Notably, the cross-coupling between electric and magnetic quantities in isotropic particles is not limited to chirality. In the chiral case (31), the off-diagonal components of the polarizability were anti-symmetric, but they may also have a symmetric part, called magnetoelectric coupling [57, 97]:
where characterizes the magnetoelectric effect of the particle. In contrast to chirality, which lacks the space-inversion P-symmetry but preserves the time-inversion T-symmetry, the magnetoelectric effect breaks both of these symmetries. Therefore, it produces a non-reciprocal response, and is sometimes called “false chirality” [58, 56, 55]. Isotropic continuous media with magnetoelectric coupling are known as “Tellegen media” [535, 440, 549, 283, 33, 229]. Remarkably, the isotropic magnetoelectric effect was first predicted for molecules by Curier and Debye more than a century ago [160, 165]. However, it still remains somewhat elusive for experimental observations.
Absorption rates, forces, and torques produced by magnetoelectric coupling can be obtained by substituting Eq. (37) into the general equations of Table 2. The main expressions have been described in [97, 239]. Importantly, while material chirality is naturally coupled to electromagnetic helicity , which is nonzero in a circularly-polarized plane wave (resulting in optical chiral dichroism as a natural measure of the material chirality), the magnetoelectric effect couples to the quantity , which vanishes in any plane wave. Nonetheless, this quantity (called “magnetoelectric density” or “reactive helicity density”) is generally nonzero in structured electromagnetic waves with locally non-orthogonal electric and magnetic fields [97, 294, 403, 230]. Therefore, the local magnetoelectric effect in interaction with small particles can be observed in structured optical fields.
@@ IV.3 Acoustic Willis coupling
As we mentioned in Section II.3, there is no acoustic counterpart of electromagnetic helicity: the chirality of a sound wave field vanishes identically due to its curl-less nature. Furthermore, it is impossible to introduce an isotropic cross-coupling, described by a single scalar parameter, like in Eqs. (31) and (37). Indeed, a cross-coupling between the scalar pressure and vector velocity fields requires a vector quantity, which implies an anisotropy. (Alternatively, vector characteristics of isotropic particles can appear in the case of spatial dispersion, i.e., a nonlocal response proportional to the wavevector of the incident field.) As a consequence, isotropic chiral particles of opposite chiralities cannot be distinguished in the linear interaction with a monochromatic sound wave field. (This can be related to the problem of Lord Kelvin’s “isotropic helicoid”: an isotropic particle that couples translation and rotation when interacting with a fluid [153].)
The cross-coupling between acoustic pressure and velocity degrees of freedom is known as Willis coupling [577, 578, 381], and it is currently attracting rapidly growing attention [389, 499, 390, 437, 374, 348, 438, 482, 483]. Willis coupling is described by the constitutive relations
where and are vector quantities, while the dipole polarizability generally becomes a tensor due to the anisotropic nature of the particle. In the reciprocal case with T-symmetry, we have [437, 482]:
This can be regarded as an acoustic counterpart of the chiral electromagnetic coupling (31). A non-reciprocal T-odd Willis coupling with is also possible [438], akin to the magnetoelectric coupling (37).
Substituting Eqs. (38) and (39) into the equations of Table 2, and assuming for simplicity a scalar dipole polarizability , we obtain the main corrections (without recoil terms) to the acoustic forces and torques induced by the Willis coupling:
The studies of Willis-coupling-induced forces and torques have started only recently [482, 483]. Evidently, these quantities depend strongly on the direction of the vector polarizability , and therefore can be attributed to anisotropic forces and torques [513]. Notably, since the acoustic recoil force terms in Eqs. (24) are aligned with the main gradient and radiation-pressure forces, the Willis-coupling force (40) can serve as the main mechanism providing a nonzero force component in the ‘third’ direction.
@@ IV.4 Generic bi-anisotropic particles
The most general small particles with anisotropy and cross-coupling between different degrees of freedom (electric and magnetic in optics or pressure and velocity in acoustics) are called bi-anisotropic particles [480]. Such particles can still be described by the same induced dipole and monopole moments, and the force and torque expressions in Table 2 are still valid. However, in this general case the electric-magnetic polarizability tensor is a complex-valued tensor, whereas the monopole-dipole polarizability in acoustics is a tensor, see Table 3. The presence of some spatial and temporal symmetries in the particle imposes some restrictions on the components of these tensors. Nonetheless, even particles of very basic shapes (such as a spiral, a cone, 2D anisotropic or chiral shapes, etc.) are characterized by a bi-anisotropic response. Explicit values of optical and acoustic forces and torques should be analysed in each specific case using the corresponding polarizability tensor and general equations in Table 2. Table 3 summarizes examples and polarizability tensors of isotropic, anisotropic, bi-isotropic, and bi-anisotropic particles in optics and acoustics.
@@ IV.5 Role of resonances
We now return to the case of isotropic particles to consider the effects of frequency dependence (dispersion) of the particle polarizabilities, forces, and torques. These are most pronounced in the vicinity of various resonances, such as the first (dipole and monopole) Mie resonances, plasmonic resonances, atomic resonances, etc. A typical resonance is characterized by the resonant frequency and the quality factor , related to the decay rate parameter as . In the vicinity of a typical resonance, the corresponding polarizability can be written in the form [509, 442, 429] (see also Appendix F)
@@
Here, for the acoustic dipole and monopole polarizabilities one should use the corresponding factors and , Eq. (20), whereas and are the decay rate coefficients related to the radiation and absorption losses [442, 314, 594, 92, 45]. The total decay rate is . The polarizability (42), as any response function in linear physical systems, obeys the Kramers–Kronig relations [544, 320].
@@
The frequency dependence of the real and imaginary parts of the polarizability (42) are shown in Fig. 10(a,b). One can see that both of these can be significantly enhanced in the vicinity of the resonant frequency. Hence, the optical or acoustic absorption rates, forces, and torques also exhibit resonant behaviour. Importantly, , and hence the gradient forces in Eqs. (22), change sign when crossing the resonant frequency, i.e., the particle is either attracted to or repelled from the high-intensity field regions, Fig. 10(c) [361]. This can be employed for various trapping applications, such as the atom trapping near a nanofiber with evanescent optical fields of two frequencies with and [319, 188, 557, 238].
@@
In turn, does not change its sign and reaches its peak value at . This provides conditions for resonant enhancement of the absorption rates (21), radiation-pressure forces in Eqs. (22) [40], and torques (23) [488], Fig. 10(d). The maximum peak value of is achieved for the negligible absorption, . This corresponds to the maximum radiation-pressure force [441].
@@
In turn, the radiation torque (26) is determined by the absorption cross-section , Eqs. (III.3). Substituting Eq. (42), we obtain
@@
Remarkably, the maximum peak value of is achieved under the so-called critical coupling conditions, when and [441, 144]. The critical coupling maximizes the amount of absorbed energy, such that the corresponding amplitude of the scattered (radiated) field vanishes [521, 462]. In some systems, this can lead to the total absorption of the incident field [594, 92, 91, 34, 45], also known as “coherent perfect absorption” [149, 561, 129]. In the context of our work, the critical coupling maximizes the optical and acoustic radiation torques [441]. Note that this single-resonance limit can be exceeded when multiple resonances are overlapped, which is called “super-absorption” [313, 462].
@@
Since the Mie resonance frequencies depend on the particle size , the resonant behaviour of the radiation forces and torques can be used for size-dependent sorting of particles [494]. It was also suggested that Helmholtz-resonator particles can provide resonant enhancement of the acoustic Willis-coupling-induced forces and torques (40) and (41) [483].
@@ IV.6 Higher multipoles
So far, we mostly considered the Rayleigh-particle regime, , which allowed us to keep the scattered fields within the lowest monopole-dipole modes. For larger Mie particles with the higher-order multipoles can play an important role in the scattered fields, and, hence, in the wave-induced forces and torques. In such case, the expressions in Table 2 are no longer applicable.
In the general case, both in optics and acoustics, the scattered field can be expanded in an infinite series of multipoles (spherical harmonics) [107, 574]. Substituting the multipole series into the general flux equations (12)–(16) and evaluating the integrals akin to Appendix C, one can calculate contributions of different multipoles to the absorption rate, force, and torque. Such approach has been implemented for optical forces [284, 624, 285, 143] and torques [568, 593], while in acoustics it is still an open challenge. A related approach, involving the multipole decompositions of both the incident and scattered fields, is known as the T-matrix formalism [399, 247, 59, 60, 240, 500, 502, 504].
Here we do not present cumbersome expressions involving higher-order multipole contributions to the radiation forces and torques, and only highlight their main features. First, the force and torque expressions are still split into the ‘mixed’ and ‘scattered’ (recoil) terms. The ‘mixed’ terms are split into independent contributions describing the coupling of the -th multipole with the incident field, whereas the recoil force terms describe the interference between the neighboring -th and -th multipoles, Fig. 11. In electromagnetism, an interference between electric and magnetic multipoles of the same order also contributes to the recoil force terms. In contrast, the higher-order contributions to the recoil optical torque include only interference of multipoles of the same nature (either electric or magnetic), and only of the same order [593, see SM]. Note that the higher-order multipole contributions to the force and torque involve the higher-order derivatives of the incident field.
Note also that one can introduce the multipole moments and the corresponding polarizabilities of the particle, similar to the dipole and monopole quantities considered above. For spherical Mie particles, the -th multipole polarizability is proportional to the -th Mie scattering coefficient [320]. For complex-shaped particles the polarizabilities can be calculated numerically [73].
The higher-multipole effects in radiation forces and torques have been studied both theoretically [302, 519, 451, 494, 490] and experimentally [40, 361, 435, 623, 624]. In particular, the multipolar decomposition is useful for the description of pulling forces [143, 140], trapping effects [330, 361], and various torque phenomena [142, 542].

### v3

@@ IV.2.1 Chiral particles
Another important case, specific to electromagnetism, is the cross-coupling between the induced electric and magnetic dipoles and the corresponding incident fields. This coupling occurs for isotropic chiral particles that lack mirror and space-inversion (P) symmetries. Such particles typically consist of chiral micro-elements (e.g., molecules) with a prevailing handedness, randomly oriented to ensure macroscopic isotropy. In this case, the induced electric and magnetic dipoles are given by [60]
@@
Substituting Eq. (33) into the general equations of Table 2, and using definitions (3)–(8), we obtain additional absorption rates, forces, and torques produced by the chirality and linear in [148, 240, 252]:
@@
These equations include both the main (‘mixed’) and recoil terms, and require further explanation.
@@
First, the chirality-dependent absorption rate (IV.2.1) describes chiral dichroism, which is central to optical sensing of chiral molecules [60, 585, 586, 275, 156, 101]. The leading term is proportional to the electromagnetic helicity density (8), which can also be associated with “electromagnetic chirality” [585, 103, 129]. The additional recoil term involves a natural complex extension of the helicity (8): , discussed in [101, 313, 440].
@@
Second, the chiral force (IV.2.1) contains three terms that parallel the gradient, radiation-pressure, and recoil forces in Eqs. (24) and (26). Specifically, the chiral gradient force is determined by the gradient of the helicity density in the incident field, whereas the chiral radiation-pressure force is proportional to the quantity , which can be called “chiral momentum” density [135, 101, 615]:
@@
where Maxwell’s equations were used. The chiral momentum density characterizes the difference between the momenta of the right-hand and left-hand circularly-polarized photons. For the helicity eigenstates , , it equals . Importantly, the chiral gradient and radiation-pressure forces in Eq. (IV.2.1) have been successfully exploited for optical sorting (separation) of chiral particles [135, 101, 130, 594, 596, 625, 328] (see Section V.5).
@@
Third, we expressed the chiral optical torque (36) using the chiral absorption cross-sections [252]
@@
As in the isotropic-particle case, Eqs. (28), the chiral torque is determined by absorption and vanishes for lossless particles. This holds for isotropic chiral spheres of any size [148], consistent with the optical theorem for chiral particles [75]. The torque (36) provides an additional mechanism of optical torque, Fig. 9(d). Even a linearly-polarized incident plane wave can induce such a torque, because an absorptive chiral particle scatters it into a field carrying non-zero angular momentum. To our knowledge, this type of torque has not yet been observed experimentally, but it should be accessible in systems with chiral metallic nano-structures [376, 238, 658].
@@
In addition to the chiral effects linear in , there are small corrections to the isotropic-particle force and torque, Eqs. (26) and (28): and . However, these corrections are insensitive to the sign of chirality and therefore cannot distinguish enantiomers; they only provide small, usually negligible corrections to the magnitudes of the isotropic recoil force and torque.
@@
Note that some works group the chiral and non-chiral force terms differently, introducing additional ‘vortex’ and ‘spin-curl’ contributions [428, 625, 356, 146]. This difference arises from not explicitly using the canonical momentum and chiral momentum (37). The framework adopted here, based on these quantities, provides a clear and physically intuitive description. A detailed comparison of equivalent formulations can be found in [252].
@@ IV.2.2 Magnetoelectric coupling
Notably, the cross-coupling between electric and magnetic quantities in isotropic particles is not limited to chirality. In the chiral case (33), the off-diagonal components of the polarizability tensor are anti-symmetric. However, they may also contain a symmetric part, corresponding to the magnetoelectric coupling [60, 101]:
where quantifies the magnetoelectric response of the particle. In contrast to chirality, which breaks space-inversion P-symmetry but preserves time-inversion T-symmetry, the magnetoelectric effect violates both. Consequently, it produces a non-reciprocal response and is sometimes referred to as “false chirality” [61, 59, 58]. Isotropic continuous media with magnetoelectric coupling are known as “Tellegen media” [590, 479, 608, 300, 36, 242]. Remarkably, the isotropic magnetoelectric effect was first predicted for molecules by Curie and Debye more than a century ago [169, 174]. However, it has remained elusive in direct experimental observations.
Absorption rates, forces, and torques produced by magnetoelectric coupling can be obtained by substituting Eq. (39) into the general equations of Table 2. The main expressions have been described in [101, 252]. Importantly, while material chirality couples naturally to the electromagnetic helicity (nonzero in a circularly-polarized plane wave and giving rise to optical chiral dichroism), the magnetoelectric effect couples to the quantity , which vanishes in any plane wave. Nonetheless, this quantity (called “magnetoelectric density” or “reactive helicity density”) is generally nonzero in structured electromagnetic waves with locally non-orthogonal electric and magnetic fields [101, 313, 440, 243]. Therefore, the local magnetoelectric response of small particles can be revealed in suitably structured optical fields.
@@ IV.3 Acoustic Willis coupling
As we mentioned in Section II.3, there is no acoustic counterpart of electromagnetic helicity: the chirality of a sound wave field vanishes identically due to its curl-less nature. Moreover, an isotropic cross-coupling, described by a single scalar parameter, as in Eqs. (33) and (39), cannot be introduced. Indeed, cross-coupling between the scalar pressure and vector velocity fields requires a vector quantity, which implies anisotropy. (Alternatively, effective vector properties may arise for isotropic particles in the presence of spatial dispersion, i.e., a nonlocal response proportional to the wavevector of the incident field.) As a consequence, isotropic chiral particles of opposite handedness cannot be distinguished in linear interactions with a monochromatic sound wave field. (This is reminiscent of Lord Kelvin’s “isotropic helicoid”: a hypothetical isotropic particle that couples translation and rotation when interacting with a fluid [162].)
The acoustic cross-coupling between the pressure and velocity degrees of freedom is known as Willis coupling [637, 638, 416], which has recently attracted intense interest [426, 552, 427, 476, 409, 378, 477, 532, 533]. The Willis coupling is described by the constitutive relations
where and are vector quantities, while the dipole polarizability generally becomes a tensor due to particle anisotropy. The reciprocal (T-symmetric) case is characterized by the following relations [476, 532]:
This can be regarded as the acoustic counterpart of the chiral electromagnetic coupling (33). A non-reciprocal T-odd Willis coupling with is also possible [477], analogous to the magnetoelectric coupling (39).
Substituting Eqs. (40) and (41) into the general equations of Table 2, and assuming for simplicity a scalar dipole polarizability , we obtain the leading corrections to the acoustic forces and torques induced by the Willis coupling (excluding recoil terms):
Studies of Willis-coupling-induced forces and torques are just emerging [532, 533]. These effects depend sensitively on the direction of the vector polarizability , and can thus be attributed to anisotropy-induced forces and torques [567]. Notably, since the acoustic recoil force terms in Eqs. (26) are aligned with the main gradient and radiation-pressure forces, the Willis-coupling force (42) can serve as the main mechanism for generating a nonzero force component in the ‘third’ direction.
@@ IV.4 Generic bi-anisotropic particles
The most generic small particles with anisotropy and cross-coupling between different degrees of freedom (electric and magnetic in optics or pressure and velocity in acoustics) are known as bi-anisotropic particles [530]. Such particles are still described by the induced dipole and monopole moments, and the general force and torque expressions in Table 2 remain applicable. However, in this case the polarizability tensors take their most general form: a complex-valued tensor for the electric–magnetic response in optics, and a tensor for the monopole–dipole response in acoustics, see Table 3. Spatial and temporal symmetries of the particle impose constraints on the tensor components, but even very simple geometries (such as spirals, cones, or 2D anisotropic and chiral shapes) exhibit a bi-anisotropic response. Explicit values of optical and acoustic forces and torques should be analyzed in each specific case using the corresponding polarizability tensor and expressions in Table 2. Table 3 summarizes representative examples and corresponding polarizability tensors of isotropic, anisotropic, bi-isotropic, and bi-anisotropic particles in optics and acoustics.
@@ IV.5 Role of resonances
We now return to the case of isotropic particles to examine the effects of frequency dependence (dispersion) in their polarizabilities, as well as in the resulting forces and torques. These effects are most pronounced in the vicinity of various resonances, such as the first (dipole and monopole) Mie resonances, plasmonic resonances, atomic resonances, etc. A typical resonance is characterized by a resonant frequency and a quality factor , which is related to the decay rate parameter by . In the vicinity of a typical resonance, the corresponding polarizability can be written in the form [562, 481, 468] (see also Appendix F)
@@
Here, for the acoustic dipole and monopole polarizabilities one should use the corresponding factors and , Eq. (22), whereas and are the decay rates associated with radiation and absorption losses, respectively [481, 335, 656, 96, 48]. The total decay rate is . The polarizability (44), like any response function in a linear physical system, must satisfy the Kramers–Kronig relations [603, 341].
@@
The frequency dependencies of the real and imaginary parts of the polarizability (44) are shown in Fig. 10(a,b). Both parts can be significantly enhanced in the vicinity of the resonant frequency. Consequently, the optical or acoustic absorption rates, forces, and torques exhibit resonant behavior. Importantly, , and hence the gradient forces in Eqs. (24), change sign when crossing the resonance. As a result, the particle can either be attracted to or repelled from high-intensity field regions, Fig. 10(c) [497, 393]. This effect can be employed for various trapping applications, such as atom trapping near a nanofiber with evanescent optical fields of two frequencies with and [340, 198, 616, 251].
@@
In turn, does not change sign and reaches its peak value at . This provides conditions for the resonant enhancement of absorption rates (23), radiation-pressure forces in Eqs. (24) [43, 365, 366], and torques (25) [540], Fig. 10(d). The maximum peak value of is achieved when absorption is negligible, . This corresponds to the maximum radiation-pressure force [480].
@@
In turn, the radiation torque (28) is determined by the absorption cross-section , Eqs. (III.3). Using Eq. (44), we obtain
@@
Remarkably, the maximum peak value is achieved under the so-called critical coupling conditions, when and [480, 152]. At critical coupling, absorption is maximized while the scattered field amplitude vanishes [576, 507]. In some systems, this can lead to the total absorption of the incident field [656, 96, 95, 37, 48], also known as “coherent perfect absorption” [157, 620, 136]. In the context of our work, the critical coupling maximizes the optical and acoustic radiation torques [480]. Note that this single-resonance limit can be surpassed when multiple resonances overlap, a phenomenon known as “super-absorption” [334, 507].
@@
Since the Mie resonance frequencies depend on the particle size , the resonant behavior of radiation forces and torques can be used for size-selective sorting of particles [547]. It has also been proposed that Helmholtz-resonator particles can provide resonant enhancement of Willis-coupling-induced acoustic forces and torques (42) and (43) [533].
@@ IV.6 Higher multipoles
So far, we have mostly considered the Rayleigh-particle regime, , where the scattered fields are well described by the lowest-order monopole and dipole modes. For larger Mie particles with , higher-order multipoles contribute significantly to the scattered fields and, hence, to the wave-induced forces and torques. In this case, the expressions in Table 2 are no longer applicable.
In the general case, both in optics and acoustics, the scattered field can be expanded into an infinite series of multipoles (spherical harmonics) [112, 634]. By substituting this multipole expansion into the general flux equations (14)–(18) and evaluating the integrals in a manner similar to Appendix C, one can calculate the contributions of different multipoles to the absorption rate, force, and torque. Such approach has been implemented for optical forces [670, 692, 302, 151] and torques [628, 655], whereas its application in acoustics remains an open challenge. A related framework, which involves the multipole decompositions of both incident and scattered fields, is known as the T-matrix formalism [436, 260, 62, 63, 253, 553, 555, 557].
Here we do not present cumbersome expressions for higher-order multipole contributions to the radiation forces and torques, but instead highlight their main features. First, the force and torque expressions are still split into ‘mixed’ and ‘scattered’ (recoil) terms. The ‘mixed’ terms contain independent contributions from the coupling of the -th multipole with the incident field, whereas the recoil force terms arise from the interference between the neighboring -th and -th multipoles, Fig. 11. In electromagnetism, interference between electric and magnetic multipoles of the same order also contributes to the recoil force terms. By contrast, the higher-order contributions to the recoil optical torque include only interference between multipoles of the same type (either electric or magnetic), and only within the same order [655, see SM]. Note that higher-order multipole contributions to forces and torques involve higher-order derivatives of the incident field.
One can also define multipole moments and the corresponding multipole polarizabilities of the particle, similarly to the dipole and monopole quantities considered earlier. For spherical Mie particles, the -th multipole polarizability is proportional to the -th Mie scattering coefficient [341]. For complex-shaped particles, the polarizabilities can be computed numerically [76].
The role of higher multipoles in radiation forces and torques has been explored both theoretically [321, 573, 494, 547, 598, 543] and experimentally [43, 393, 600, 474, 691, 692]. In particular, the multipolar decomposition has proven useful in describing chiral forces [266, 536, 134, 688, 454], pulling and lateral forces [151, 147, 429], trapping effects [353, 393, 600], as well as torque engineering in the linear [150, 601] and nonlinear regimes [597].